▾▴ Toggle Texts
▾▴ Toggle Images
◱ Full Size Pics
⊻ Bottom of Page
- (inserted: 17.06.2022 06:19)
Sorry for the delayed report.Sorry for the delayed report. It had been very hot here and we were a bit fried yesterday. On Wednesday we logged the bore hole and installed the strainmeter at 115.5 m that afternoon. About as perfect of an install as you can get. We installed the seismometer today at 110 m and the screen between 82 – 91 m to sample a water bearing fracture at 84 m. Tomorrow we will complete the downhole work at TSM6.
- (inserted: 16.06.2022 07:52)
We drilled to 118 m today and decided to stop and flush the borehole.We drilled to 118 m today and decided to stop and flush the borehole. All polymer was removed but the water was still cloudy. Despite this we stopped cleaning the borehole. The drillers tripped out without issues. The water sample taken at the end of flushing had cleared up significantly by the end of the day.
- (inserted: 08.06.2022 05:45)
Reached 27 m today.Reached 27 m today. Driller set conductor casing to 8 m and then drilled for the cased section. The outer casing is planned to a total depth of 79 m. Drilling times varied between 40 min and 83 min per 3m drill rod as the formation hardness fluctuated. Very few fracture zones so far.
- (inserted: 07.06.2022 10:03)
The drillers completed their 4-day break and returned today.The drillers completed their 4-day break and returned today. They finished setting up and are ready to drill. Unfortunately, the water truck had mechanical issues and didn’t arrive until this evening. Drilling will commence tomorrow morning.
- (inserted: 01.06.2022 10:42)
Today the drillers began their mobilisation from site5 to site 6.Today the drillers began their mobilisation from site5 to site 6. The GNSS borehole adapter was installed at site 5 and the water level in the sampling pipe is at 6.14 m. We trenched and buried the surface seismometer at TSM4 today.
- (inserted: 31.05.2022 07:47)
We completed the down hole work at site 5 today.We completed the down hole work at site 5 today. On Saturday we installed the seismometer and pore pressure pipe infrastructure. We touched the grout at 112.5 m and then pulled the seismometer up to 111 m to align the screen with a fracture zone at 95 m. Today we sandpacked the screen section, put in a bentonite cap and cemented to 2 m below surface. The drillers also drilled and cased a 12 m hole for a broadband instrument.
- (inserted: 30.05.2022 06:29)
We successfully installed the BSM (strainmeter) at 116.1m on a very warm day today (30°C).We successfully installed the BSM (strainmeter) at 116.1m on a very warm day today (30°C). For this we used gravel to bring up the bottom by 80 cm. It appears some material may have fallen to the bottom of the bore hole from the fracture at 95m bringing the hole up 30-40 cm. We are still well within our target install zone.
- (inserted: 25.05.2022 18:45)
We reached a target depth of 118m.We reached a target depth of 118m. We originally planned on a TD of 120+ m but decided the drilling indicated good rock and we did not want to push our luck. The drillers flushed the cuttings and polymer out of the bore hole and tripped out. After tripping out they filled the borehole with clean H2O. We will log tomorrow after letting the temperature in the borehole stabilize and to allow any material in the water column to settle out.
- (inserted: 24.05.2022 06:50)
Today we drilled down to 82.4 m and managed to case the hole.Today we drilled down to 82.4 m and managed to case the hole. In the next couple of days, we are going to drill down to at least 120 m, plan to flush the bore hole, trip out and log on Thursday.
- (inserted: 23.05.2022 07:26)
The drill rig and equipment moved to site 5The drill rig and equipment moved to site 5 yesterday, the drill team spudded in this morning and set conductor casing. We drilled to 18m today. Casing target is 79 m. At site 4 the GPS adapter was welded to the wellhead and we measured the water in the pore pressure pipe at 17.4m.
- (inserted: 18.05.2022 15:00)
We completed the down hole work at site 4We completed the down hole work at site 4 today by putting in a sand pack between 81-91m, a bentonite cap between 80-81m, and cementing to the surface. The drilling pad at TSM5 will be finished tomorrow and hopefully we will spud in tomorrow afternoon.
- (inserted: 17.05.2022 05:36)
We logged the well on Friday installed the strainmeterWe logged the well on Friday, installed the strainmeter at 100.3m on Saturday and the seismometer on Monday at 95.1. The screen is at 85 - 91m. It is panned to sand in the screen section, put in a bentonite cap, and cement up the borehole on Tuesday. We should able to mobilize over to site 5 by Thursday.
- (inserted: 12.05.2022 18:39)
Today we made it our target depth of 100.5Today, we made it our target depth of 100.5 m and decided not to go any deeper because there appears to be an install zone between 98 -100 m. We will log tomorrow after cleaning out the borehole and put together the strainmeter, which we moved over to the site today.
- (inserted: 09.05.2022 14:30)
Today the science team visited sites 5 and 6Today the science team visited sites 5 and 6 to discuss the location of the borehole and pads with the land owners (site 6) and the municipality (site 5). Pads and mud pits should be in before the rig is moved. We also measured the water level at site 3 (14.8 m).
- (inserted: 05.05.2022 09:16)
Good day of drilling today with up to 3mGood day of drilling today with up to 3m every 30 min by the end of the day, reaching 39 m. This was despite hitting some clay layers early on that slowed down the drilling due to slowing down the bit rotation
- (inserted: 05.05.2022 09:13)
Good progress today reaching a depth of 8.5 mGood progress today, reaching a depth of 8.5 m, so the drillers can install the conductor casing to at least 10m tomorrow. The soil and upper section of the rock is clay rich and should form a good seal with the conductor casing. Due to proximity to active CO2 out gassing we are being very cautious and continuously monitor for CO3, H2S, and O2 levels. The drillers are using a diverter in case we hit an over pressurized section of rock. We used a similar system in Yellowstone NP.
- (inserted: 24.11.2021 07:45)
We completed all downhole work at TSM03 today includingWe completed all downhole work at TSM03 today including the sand and bentonite pack around the screen (60-69 m) and cementing to the surface. The enclosure was built at TSM02 and we got the communication system working at TSM01. It was turned off while we wait on power. Tomorrow we will prepare TSM03 for its enclosure pads and install solar and communications at TSM02.
- (inserted: 24.11.2021 07:43)
We installed the BSM at 78.6m at TSM03 today.We installed the BSM at 78.6m at TSM03 today. We had reached 80 m on Friday and logged the borehole on Saturday. The logged depth was 79.4m. Logs showed very consistent rock but it had slow zones on the FWS at 78.6 and 76.3m. We lifted the hole to 78.6 with gravel. The install went very smoothly with no issues. Tomorrow we will install the seismometer and pore pressure tube.
- (inserted: 12.11.2021 07:54)
We drilled down to 61m today.We drilled down to 61m today. Yesterday the drillers drilled and cased down to 32.4m. The inclination was tested this morning at 31.4m and was 0.6 degrees from vertical. The plan is to reach 80 m tomorrow and bring the borhole seismometer SM up to site; logging on Saturday morning.
- (inserted: 10.11.2021 09:30)
We had some bad luck at TSM03 today.We had some bad luck at TSM03 today. The drillers had reached 41m from the top of surface casing today. However, while installing casing today a thread failed causing four 3m sections to be stuck in the hole. The drillers were able to retrieve three sections. Unfortunately, the fourth section was stuck. The safest, most time, and cost-effective decisions was to abandon the borehole. The rig was moved to 2 m to the north and we managed to drill 1 m before the day ended.
- (inserted: 09.11.2021 07:37)
The drillers made good progress at site 3 (TSM03) today.The drillers made good progress at site 3 (TSM03) today. They set 9 m of surface casing and reached 31 m total depth. Plan is to drill to 40 m and begin casing tomorrow.
- (inserted: 07.11.2021 15:06)
The drillers moved from site 2 (TSM02) to site 3 (TSM03) today.The drillers moved from site 2 (TSM02) to site 3 (TSM03) today. They were very fortunate that the rain forecast was wrong today and equipment was moved without any issues. Work will be resumed on Monday Nov 8th.
- (inserted: 05.11.2021 17:23)
Today, the team installed the seismometer, fiber optic cable, and pore pressure tube at site 2...Today, the team installed the seismometer, fiber optic cable, and pore pressure tube at site 2 (TSM02). It was a great team effort: The borehole seismometer (BSM) was cemented in at 152.5 m depth and the top of the BSM grout was at 153m. the team hopes to complete the down hole work at TSM02 on Thursday and begin mobilization at TSM03. The BSM is running on solar power. The screen is installed at 139.5 to 148.5 m depth.
- (inserted: 03.11.2021 18:11)
Today, we had a successful strainmeter installation at TSM02.Today, we had a successful strainmeter installation at TSM02. It was installed at 158.1m. The well was logged on Tuesday evening and luckily the best rock was near the bottom of the borehole.
- (inserted: 03.11.2021 08:31)
The drillers reached their target depth on Saturday October 30th.The drillers reached their target depth on Saturday October 30th. The final depth was 160.9m from the top of the casing. The drillers will return to the site on Tuesday to flush the borehole for logging and hopefully we can install the strainmeter on Wednesday.
- (inserted: 28.10.2021 07:03)
The drillers continued to have issues with their drillThe drillers continued to have issues with their drill collars but reached casing depth in the end and cased down to 96.8 m today. Tomorrow they will cement the casing and plan to continue drilling on Friday.
- (inserted: 26.10.2021 05:36)
On October 21st one of the drill rods threadsOn October 21st one of the drill rods’ threads failed at the top of the rod leaving 12 m of drill rod in the borehole. Fishing attempts were unsuccessful until the weekend. The drillers received a new fishing tool today and pulled out the lost rod. They reached 61 m today. We also worked on the uphole infrastructure at TSM01, which should be completed by the end of the week depending on weather.
- (inserted: 21.10.2021 06:36)
Today we began drilling the second STAR borehole afterToday, we began drilling the second STAR borehole after the drillers had moved on from TSM01 on Tuesday after a well-deserved break. They set surface casing to six meters and drilled to one meter. Work continues at TSM01 to complete the uphole installation. The strainmeter at TSM01 has been online since it was installed and recorded a 3.7 Mw approximately 70 km to the south east of the site on Monday.
- (inserted: 14.10.2021 17:00)
Today marked the last day of major downhole work at the first drill site!Today marked the last day of major downhole work at the first drill site! Sand was poured downhole around the screened section of PVC to allow equilibration with the water level. Once the water level stabilized, the pore pressure sensor was installed. A layer of bentonite was laid over the sand, and the remainder of the hole was cemented. The strainmeter is hooked up to a temporary solar panel setup for the long weekend. So far the data looks good.
- (inserted: 13.10.2021 17:41)
After the grout for the strainmeter cured for almostAfter the grout for the strainmeter cured for almost a day, we prepared for strainmeter and fiber optic cable installation today. The seismometer was lowered downhole on a pvc pipe with the cable attached and cemented. Both, the instrument and the bottom of the fiber optic cable are in place! The depth to the top of the grout is ~127.4 m. All gauges of the strainmeter are responding to the grout curing signal.
- (inserted: 12.10.2021 07:27)
The drillers reached 133 m today, flushed the hole until no polymer remained, and tripped out.The drillers reached 133 m today, flushed the hole until no polymer remained, and tripped out. In the afternoon, the loggers ran temperature, natural gamma, a three-arm caliper, acoustic and optical televiewers, and full wave sonic. Meanwhile, we prepped the strainmeter for a night of testing and checked the seismometer functionality
- (inserted: 12.10.2021 07:24)
The drilling went well today!The drilling went well today! After reassessing the drillers confirmed that they were at 99 m depth at the time of casing. Today, drilling finished to 127 m. There were a few minor fractures downhole. On the geochemistry side of things, we took down the spectrometer. The photo shows Adriano (INGV) measuring the mud pit pH: it was basic like expected.
- (inserted: 12.10.2021 07:22)
Today, the drillers cemented outside of the casing again.Today, the drillers cemented outside of the casing again. They had return flow of cement up to the surface, so the large fracture zone was passed and the cement is currently drying. We sent the inclinometer downhole to assess the overall deviation - a maximum of 5 degrees was considered allowable, and the measurement read 0.6 degrees! Tomorrow the plan is to drill beyond the casing.
- (inserted: 12.10.2021 07:10)
Today, we tested the logging gear at the storage site and unbox one of the strainmeters.Today, we tested the logging gear at the storage site and unbox one of the strainmeters. Unfortunately, the stepdown transformer for the logging gear failed during the test. A new one will arrive on Saturday. The casing for the upper 99 m of the borehole arrived this afternoon.
- (inserted: 05.10.2021 18:17)
It is likely that we transitioned into the Maiolica yesterday at around 70 m depth when circulation...It is likely that we transitioned into the Maiolica yesterday at around 70 m depth when circulation was lost. Since this is the formation the instruments will be installed within, the drill team stopped drilling at a depth of 99 m and tripped out today. They plan on receiving the casing for the hole tomorrow.
- (inserted: 05.10.2021 07:35)
The gas collection system was setup and drilling was resumed this morning.The gas collection system was setup and drilling was resumed this morning. Near 73 m depth, there was no mudflow return due to the presence of fractures. It returned at around 80 m, however, with continued mud loss. With four water truck deliveries, the day ended at 92.5 m depth.
- (inserted: 03.10.2021 15:02)
The geochemistry team arrived at the site and resumed gas collection.The geochemistry team arrived at the site and resumed gas collection. It was noticed that the polymer in the mud pit is likely contributing CO2 and will need to be analzyed in closer detail to understand effects on the concentrations. Further, issues with the mud flow system were solved. By the end of the work day the drill had reached 67 m depth - a great way to end the week!
- (inserted: 01.10.2021 16:24)
Drilling continues at 10 m depth drilling reached theDrilling continues: at 10 m depth drilling reached the consolidated rock and the team was able to set up the equipment for online gas analysis. Every 3 meters samples of the drill cuttings were collected and labeled for future use. Today, the final depth was 23.3 m.
- (inserted: 29.09.2021 19:44)
Today after having put the drill rig in place, we started the proper drilling of the fist borehole.Today after having put the drill rig in place, we started the proper drilling of the fist borehole. So far we drilled 9 meters and we have been able to put the conductor casing in place as well.
- (inserted: 05.05.2022 09:09)
The 2nd operational phase of STAR started today LauroThe 2nd operational phase of STAR started today: Lauro Chiaraluce and Wade Johnson arrived on site and the excavation crew was building the pad and reinforcing the road to the site. The transport of the drill rig is slightly delayed.